Investing $2,000,000 in TQM's Channel Support Systems initiative will at a minimum increase demand for your products 3.0% in this and in all future rounds. (Refer to the TQM Initiative worksheet in the CompXM Decisions menu.) Looking at the Round 0 Inquirer for Andrews, last year's sales were $163,085,264. Assuming similar sales next year, the 3.0% increase in demand will provide $4,892,558 of additional revenue.   With the overall contribution margin of 34.1%, after direct costs this revenue will add $1,668,362 to the bottom line. For simplicity, assume that the demand increase and margins will remain at last year's levels. How long will it take to achieve payback on the initial $2,000,000 TQM investment, rounded to the nearest month?
Select: 1
TQM investment will not have a significant financial impact
5 months
14 months
10 months
Answer #1

Correct answer is option b.14 months

Last year’s sales = $163,085,264

As per the given data next year sales is increased by 3.0%. So, 3.0% of last year sales

= 0.03 * $163,085,264

= 4,892,558

It has given that Revenue added to the bottom line = 34.1% of increased demand

= 0.341 * 4,892,558= 1668362

Payback on the initial $2,000,000 TQM investment= (Investment in TQM / Revenue added to the bottom line) * 12

= ($2,000,000 / $1668362)*12 = 14.39 = 14 Months